# When Location Shapes Choice:

### v3

@@ 1 Introduction
An underlying assumption in most discrete choice models is that a customer’s choice depends only on his/her preferences over the assortment of products. In practice, when a customer browses a store - whether physical or online - he/she typically encounters only a subset of the products on display. Customers chooses based on their preferences over the products that they actually see, and the likelihood of observing a given product depends on the location of this product in the store. Some locations may have high visibility to customers and a product placed in such a location would have a higher likelihood of being seen by the customer (Larson et al. (2005), Hui et al. (2009), Agarwal et al. (2011), Ghose and Yang (2009), Ursu (2018)). In such circumstances, customer choice behavior is influenced by the placement of products at different locations as well as the overall assortment of products on display.
@@
We model the customer behavior as a two-stage process: first, a customer entering the store visits a subset of display locations according to a browsing distribution; second, the customer chooses from the products placed at the locations that they visited according to a discrete choice model. The browsing distribution is the probability distribution of the set of visited locations. Note that customers may take different paths through the store (for example, see Figure 1). A customer’s path may depend on the layout of the store and the customer’s patience level and capacity for browsing but we assume that it does not depend on the product placement, i.e., the browsing distribution is independent of the placement of products. We also assume that in the second stage, the customer’s choice depends only on the set of products that they saw in the first stage and not on where these products were displayed, i.e., the discrete choice model is also independent of the placement of products. Later, we generalize some of our results to settings where the choice model depends on the placement.
The objective is to select an assortment of products and determine their placement in order to maximize the expected revenue from a customer visit. Each location holds at most one product but a given product may be placed at multiple display locations. Our model also captures settings where multiple products may be placed at each location. To capture the different types of choice behavior and browsing patterns that may arise in the diverse array of applications of this problem, we focus on designing a general algorithm that can select and place the products optimally for any browsing distribution and almost any choice model. We call this the Placement problem.
@@
In the setting where customers visit all locations with certainty (a trivial browsing distribution), the placement problem is equivalent to the classical assortment optimization problem with a cardinality constraint due to the limited number of display locations. For a general choice model such as the Random Utility Model (RUM), this problem is computationally intractable and no non-trivial approximation guarantees are possible (Aouad et al. (2018)). However, for many special cases of RUM that capture single product purchase behavior, such as the Multinomial Logit (MNL) choice model, Nested Logit (NL) choice model, or the Markov choice model, there exist efficient algorithms or approximation schemes for finding the optimal or near-optimal assortment under a cardinality constraint (see Section 1.2). We show that an assortment that is optimal when customers visit all locations may not be a good assortment of products for other (non-trivial) browsing distributions. More broadly, while the literature on constrained assortment optimization is extensive, it is not immediately apparent if its algorithmic results can be used for the Placement problem.
@@ 1.1 Main contributions
Our goal is to design an efficient algorithm for a general setting of the Placement problem. Prior work on assortment optimization has demonstrated the benefits of using the specific characteristics of a choice model in order to design provably good and efficient algorithms. We focus on developing a general algorithmic framework that can generate a provably good placement solution for any given choice model and browsing distribution by utilizing an algorithm for assortment optimization in the same choice model. Such a framework allows us to transform a customized assortment optimization algorithms for a given choice model into an approximation algorithm for Placement with the same choice model.
Placement for Maximizing Revenue: Our main technical contribution is a general approximation algorithm for the Placement problem when products have fixed prices and the objective is to maximize expected revenue. Our result holds for a very general family of choice models, including models for multi-purchase behavior. Specifically, we only require the choice model to satisfy the weak-rationality and monotonicity properties, a condition met by a large class of single-purchase choice models, including the RUM, and several multi-purchase choice models. Given a choice model that satisfies this condition, any browsing distribution, and an oracle (black box algorithm) that provides an $\alpha$-approximate solution to the cardinality constrained assortment optimization problem under the given choice model, we give an algorithm that achieves an approximation ratio of $\frac{\Theta(\alpha)}{\log\rho}$ for the Placement problem, where $\rho$ is a problem parameter that depends on the variation in the number of locations visited by different customers. We have $\log\rho\approx 1$ if all the customers visit nearly the same number of locations, regardless of the total number of display locations. In the worst case, $\rho=m+1$, where $m$ is the total number of display locations, so we get $\frac{\Theta(\alpha)}{\log m}$-approximation for the Placement problem.
@@
Placement for Maximizing Sales: Sales maximization is a special case of revenue maximization. Specifically, one can find a placement that optimizes expected sales by solving an instance of the revenue maximization problem with identical prices for all the products. When prices are identical and the choice model is (single-purchase) RUM, the problem of cardinality constrained assortment optimization reduces to an instance of cardinality constrained submodular maximization for a general class of choice models (Berbeglia and Joret (2020)). We prove that the more general Placement problem for products with uniform/identical prices is equivalent to maximizing a monotone submodular function under a matroid constraint. Submodular maximization subject to a matroid constraint is a well-studied problem that has a $(1-1/e)$-approximation algorithm (Calinescu et al. (2011)). This gives us a $(1-1/e)$-approximate algorithm for Placement with uniform prices. We also show that this is the best possible approximation guarantee for any polynomial time algorithm (unless P=NP). Our hardness result is based on the hardness of finding the optimal cardinality constrained assortment for a Mixture of MNL choice model when products have identical (or uniform) prices.
@@ 1.2 Related Literature
The impact of product placement has been studied in a variety of fields including operations research, marketing, and computer science. Recall that in Placement we consider a consider-then-choose model where the probability that a product is in the consideration set depends on its placement. We assume that the browsing distribution and the choice model are known and do not depend on the placement. Our goal is to find the revenue optimizing assortment and product placement. This problem has connections to a variety of formulations considered in the literature.
@@
Product Framing and Display Optimization: The literature on product framing and display optimization is perhaps the most closely related to our work (Aouad and Segev (2021), Davis et al. (2015), Gallego et al. (2020), Asadpour et al. (2023)). This literature considers settings where customers browse through the display locations in a fixed direction, such as from top to bottom in a list of ads on an online search platform, and then choose at most one product from the set of products that they see. As we discuss in Section 2, this corresponds to a special case of our setting where the browsing distribution is supported on nested sets of locations. We call this problem Placement on a Line. To the best of our knowledge, finding efficient algorithms for general browsing distributions remained open prior to our work. Most of the literature on Placement on a Line assumes that the browsing distribution and choice model are independent of the placement of products. There are exceptions, such as Gallego et al. (2020) and Aouad et al. (2022), that we discuss in more detail below.
@@
Davis et al. (2015) introduce the problem of assortment over time where retailers build their assortment incrementally. Their setting is a special case of Placement on a Line where the browsing distribution is uniform. They give a $0.5\alpha$-approximation for the problem for any monotone choice model that admits an $\alpha$-approximation to the cardinality constrained assortment optimization problem. Gallego et al. (2020) consider a generalization of this setting where the browsing distribution is non-uniform. They give a $6\alpha/\pi^{2}$-approximation algorithm when the browsing distribution satisfies the new better than used in expectation (NBUE) property. They also develop novel approximation algorithms for settings where the choice model either depends on the product placement or depends on the number of locations visited by the customer. Aouad and Segev (2021) consider Placement on a Line with MNL choice model and non-uniform browsing distribution. They give a polynomial time approximation scheme (PTAS) based on approximate dynamic programming. A distinctive feature of their model is that a product cannot be displayed at multiple locations in the list (and no location can be left empty) and this makes the optimization problem more challenging. Feldman and Jiang (2021) consider a model with additional “desirability” constraints on the products displayed at each level and give a PTAS. Asadpour et al. (2023) consider Placement on a Line with additional constraints and develop a $(1-1/e)$-approximation for the setting where all products have the same price.
@@
Beyond Placement on a Line, Aouad et al. (2022) considered the problem of placing art in a gallery and proposed a Pathway-MNL (P-MNL) model that captures the browsing behavior of customers as a Markovian random walk over the various locations. In their model: (1) visitors are not necessarily making a choice between substitutes but are instead visiting several artworks before leaving the gallery and the goal is to engineer visitors’ experience, (2) each product (artwork) can be displayed at no more than one location, and (3) the visitor’s browsing distribution can depend on the placement of artwork. They show that the problem of optimizing the placement of artwork in the gallery is NP hard. They propose various heuristics to solve the problem and conduct numerical experiments to demonstrate the efficacy of these heuristics but do not analyze the performance of these heuristics theoretically. In our model, the goal is to engineer customer choice and we assume that the browsing distribution is independent of the placement decision but is otherwise arbitrary (and not necessarily Markovian) and we consider a general choice model (not just MNL). We also allow a product to be displayed at multiple locations (in line with retail practice) and focus on developing efficient algorithms with theoretical performance guarantees.
@@
One common approach is to consider a choice model over an expanded universe that consists of all (product, location) pairs. The revenue optimal placement in this model is the optimal solution to an assortment optimization problem with constraints ensuring that every location has at most one product (Abeliuk et al. (2016)). This constraint can be expressed using a partition matroid or a system of totally unimodular (TU) constraints. For MNL choice model, Avadhanula et al. (2016), Sumida et al. (2021) give a polynomial size LP formulation to solve the TU constrained assortment optimization problem. For the Markov choice model (Blanchet et al. (2016)), Udwani (2023) gave a $\frac{1}{4}$-approximation algorithm for assortment optimization subject to a matroid constraint and Désir et al. (2020) showed that assortment optimization with TU constraints is NP hard to approximate with a factor of $O(n^{1/2-\epsilon})$ for any fixed $\epsilon>0$.
@@
Sequential Search: Our model also bears some resemblance to discrete choice models based on sequential search (Weitzman (1978)). In sequential search, the customer initially does not know the value of the product at any location and searches for the best alternative by examining the locations one by one. There is a cost for examining each location and the search process induces a placement dependent browsing distribution over display locations because the customer’s decision to stop searching may depend on the set of products that they have seen. The setting considered by Derakhshan et al. (2022) is perhaps the most closely related to our work. They consider a product placement problem with sequential search where customers examine the display locations in a fixed order (similar to Placement on a Line) to form a consideration set and then choose at most one product from this set according to a Mixture of MNL choice model. They give novel approximation algorithms and polynomial time approximation schemes for various types of objectives. We refer to Wang and Sahin (2018), Kosilova (2023), Ursu et al. (2023) for more detailed reviews of the literature on assortment optimization and discrete choice models based on sequential search.
Assortment Optimization: Recall that the problem of cardinality constrained assortment optimization is a special case of our setting where customers see all products on display. The placement of products does not influence customer choice but the finite number of display locations induces a cardinality constraint. There is a long line of work on assortment optimization for single-purchase choice models. Rusmevichientong et al. (2010) gave a polynomial time algorithm for the problem for MNL choice model. For the Nested Logit choice model, Gallego and Topaloglu (2014) showed that the problem can be solved in polynomial time using linear programming. For the Mixed MNL choice model with a constant number of mixtures, there is a fully polynomial time approximation scheme (FPTAS) (Mittal and Schulz (2013), Désir et al. (2022)). Désir et al. (2020) gave a $0.5-\epsilon$ approximation for the problem (with algorithm runtime proportional to $\frac{1}{\epsilon}$) in the Markov choice model and also showed that the problem is APX-hard. Cardinality constrained assortment optimization has also been studied for various consider-then-choose choice models with placement independent consideration sets. We refer to Aouad et al. (2019, 2021) for a detailed review of these settings.
@@
Shelf-Space Allocation: There is a significant body of work on the problem of allocating shelf-space in physical stores that sell a variety of (not necessarily substitutable) products (Flamand et al. (2023, 2016), Hariga et al. (2007), Flamand et al. (2018)). To the best of our knowledge, the model and methodologies considered in this related literature and our paper are complementary. The common solution methodology in this literature is to formulate a mathematical program (typically a non-linear mixed integer program) that incorporates a variety of important practical considerations in store wide allocation of shelf-space for different product categories that are not substitutable. We refer to Flamand et al. (2023) for a more detailed review of work on shelf-space allocation.

### v4

@@
Strategic product placement can have a strong influence on customer purchase behavior in physical stores as well as online platforms. Motivated by this, we consider the problem of optimizing the placement of substitutable products in designated display locations to maximize the expected revenue of the seller. We model the customer behavior as a two-stage process: first, the customer visits a subset of display locations according to a browsing distribution; second, the customer chooses at most one product from the displayed products at those locations according to a choice model. Our goal is to design a general algorithm that can select and place the products optimally for any browsing distribution and choice model, and we call this the Placement problem. We give a randomized algorithm that utilizes an $\alpha$-approximate algorithm for cardinality constrained assortment optimization and outputs a $\frac{\Theta(\alpha)}{\log m}$-approximate solution (in expectation) for Placement with $m$ display locations, i.e., our algorithm outputs a solution with value at least $\frac{\Omega(\alpha)}{\log m}$ factor of the optimal and this is tight in the worst case. We also give algorithms with stronger guarantees in some special cases. In particular, we give an efficient deterministic $\frac{\Omega(1)}{\log m}$-approximation algorithm for the Markov choice model, and a tight (11/e)(1-1/e)-approximation algorithm for the problem when products have identical prices.
[…]
@@ 1 Introduction
An underlying assumption in most discrete choice models is that a customer’s choice depends only on his/her preferences over the assortment of products. In practice, when a customer browses a store - whether physical or online - he/she typically encounters only a subset of the products on display. Customers choose based on their preferences over the products that they actually see, and the likelihood of observing a given product depends on the location of this product in the store. High-visibility placements significantly boost a product’s chances of being seen by customers in both physical stores (Larson et al. 2005) and online environments such as online advertising (Ghose and Yang 2009, Agarwal et al. 2011) and e-commerce (Ursu 2018). In general, customer choice behavior is influenced by the placement of products at different locations as well as the overall assortment of products on display.
@@
We model the customer behavior as a two-stage process: first, a customer entering the store visits a subset of display locations according to a browsing distribution; second, the customer chooses from the products placed at the locations that they visited according to a discrete choice model. We refer to the probability distribution of the set of visited locations as the browsing distribution. Note that customers may take different paths through the store (for example, see Figure 1). A customer’s path may depend on the layout of the store and the customer’s patience level and capacity for browsing but we assume that it does not depend on the product placement, i.e., the browsing distribution is independent of the placement of products. We also assume that in the second stage, the customer’s choice depends only on the set of products that they saw in the first stage and not on where these products were displayed, i.e., the discrete choice model is also independent of the placement of products. Later, we generalize some of our results to settings where the choice model depends on the placement.
The objective is to select an assortment of products and determine their placement in order to maximize the expected revenue from a customer visit. Each location holds at most one product but a given product may be placed at multiple display locations. Our model also captures settings where multiple products may be placed at each location, as well as, the setting where locations cannot be left empty. To capture the different types of choice behavior and browsing patterns that may arise in the diverse array of applications of this problem, we focus on designing a general algorithm that can select and place the products optimally for any browsing distribution and almost any choice model. We call this the Placement problem.
@@
In the setting where customers visit all locations with certainty (a trivial browsing distribution), the placement problem is equivalent to the classical assortment optimization problem with a cardinality constraint due to the limited number of display locations. For a general Random Utility Maximization (RUM)-based choice model (Block and Marschak 1959), this problem is computationally intractable and no non-trivial approximation guarantees are possible (Aouad et al. (2018)). However, for many RUM-based single-purchase choice models, such as the Multinomial Logit (MNL) choice model, Nested Logit (NL) choice model, or the Markov choice model, there exist efficient algorithms or approximation schemes for finding the optimal or near-optimal assortment under a cardinality constraint (see Section 1.2). We show that an assortment that is optimal when customers visit all locations may not be a good assortment of products for other (non-trivial) browsing distributions. More broadly, while the literature on constrained assortment optimization is extensive, it is not immediately apparent if its algorithmic results can be used for the Placement problem.
@@ 1.1 Main contributions
Our goal is to design an efficient algorithm for a general setting of the Placement problem. Prior work on assortment optimization has demonstrated the benefits of using the specific characteristics of a choice model in order to design provably good and efficient algorithms. We focus on developing a general algorithmic framework that can generate a provably good placement solution for any given choice model and browsing distribution by utilizing an algorithm for assortment optimization in the same choice model. Such a framework allows us to transform a customized assortment optimization algorithm for a given choice model into an approximation algorithm for Placement with the same choice model.
Placement for Maximizing Revenue: Our main technical contribution is a general approximation algorithm for the Placement problem when products have fixed prices and the objective is to maximize expected revenue. Our result holds for a very general family of choice models, including models for multi-purchase behavior. Specifically, we only require the choice model to satisfy the weak-rationality and monotonicity properties, a condition met by a large class of single-purchase choice models, including for any RUM-based choice model, and several multi-purchase choice models. Given a choice model that satisfies this condition, any browsing distribution, and an oracle (black box algorithm) that provides an $\alpha$-approximate solution to the cardinality constrained assortment optimization problem under the given choice model, we give an algorithm that achieves an approximation ratio of $\frac{\Theta(\alpha)}{\log\rho}$ for the Placement problem, where $\rho$ is a problem parameter that depends on the variation in the number of locations visited by different customers. We have $\log\rho\approx 1$ if all the customers visit nearly the same number of locations, regardless of the total number of display locations. In the worst case, $\rho=m+1$, where $m$ is the total number of display locations, so we get $\frac{\Theta(\alpha)}{\log m}$-approximation for the Placement problem.
@@
Placement for Maximizing Sales: Sales maximization is a special case of revenue maximization. Specifically, one can find a placement that optimizes expected sales by solving an instance of the revenue maximization problem with identical prices for all the products. When prices are identical and the (single-purchase) choice model fits within the RUM framework, the problem of cardinality constrained assortment optimization reduces to an instance of cardinality constrained submodular maximization for a general class of choice models (Berbeglia and Joret (2020)). We prove that the more general Placement problem for products with uniform/identical prices is equivalent to maximizing a monotone submodular function under a matroid constraint. Submodular maximization subject to a matroid constraint is a well-studied problem that has a (11/e)(1-1/e)-approximation algorithm (Calinescu et al. (2011)). This gives us a (11/e)(1-1/e)-approximate algorithm for Placement with uniform prices. We also show that this is the best possible approximation guarantee for any polynomial time algorithm (unless P=NP). Our hardness result is based on the hardness of finding the optimal cardinality constrained assortment for a Mixture of MNL choice model when products have identical (or uniform) prices.
@@ 1.2 Related Literature
The impact of product placement has been studied in a variety of fields including operations research, marketing, and computer science. Recall that the two-stage process in Placement is inspired by consider-then-choose choice models (Aouad et al. (2021), Jagabathula et al. (2024)), where the customer first shortlists a subset of the assortment or products and then chooses from the subset according to a discrete choice model. The main difference is that the probability that a product is in the consideration set depends on its placement and in this way our model endogenizes the consideration set of customers. We assume that the browsing distribution and the choice model are known and do not depend on the placement. Our goal is to find the revenue optimizing assortment and product placement. This problem has connections to a variety of formulations considered in the literature.
@@
Product Framing and Display Optimization: The literature on product framing and display optimization is perhaps the most closely related to our work. This literature considers settings where customers browse through the display locations in a fixed direction, such as from top to bottom in a list of ads on an online search platform, and then choose at most one product from the set of products that they see. As we discuss in Section 2, this corresponds to a special case of our setting where the browsing distribution is supported on nested sets of locations. We call this problem Placement on a Line. To the best of our knowledge, finding efficient algorithms for general browsing distributions remained open prior to our work. Most of the literature on Placement on a Line assumes that the browsing distribution and choice model are independent of the placement of products. There are exceptions, such as Gallego et al. (2020) and Aouad et al. (2025a), that we discuss in more detail below.
@@
Davis et al. (2015) introduce the problem of assortment over time where retailers build their assortment incrementally. Their setting is a special case of Placement on a Line where the browsing distribution is uniform. They give a $0.5\alpha$-approximation for the problem for any monotone choice model that admits an $\alpha$-approximation to the cardinality constrained assortment optimization problem. Gallego et al. (2020) consider a generalization of this setting where the browsing distribution is non-uniform. They give a $6\alpha/\pi^{2}$-approximation algorithm when the browsing distribution satisfies the new better than used in expectation (NBUE) property. They also develop novel approximation algorithms for settings where the choice model either depends on the product placement or depends on the number of locations visited by the customer. Aouad and Segev (2021) consider Placement on a Line with MNL choice model and non-uniform browsing distribution. They give a polynomial time approximation scheme (PTAS) based on approximate dynamic programming. A distinctive feature of their model is that a product cannot be displayed at multiple locations in the list and no location can be left empty. This makes the optimization problem more challenging. Feldman and Jiang (2023) consider a model with additional “desirability” constraints on the products displayed at each level and give a PTAS. Asadpour et al. (2023) consider Placement on a Line with additional constraints and develop a (11/e)(1-1/e)-approximation for the setting where all products have the same price.
@@
Beyond Placement on a Line, Aouad et al. (2025a) considered the problem of placing art in a gallery and proposed a Pathway-MNL (P-MNL) model that captures the browsing behavior of customers as a Markovian random walk over the various locations. In their model: (1) visitors are not necessarily making a choice between substitutes but are instead visiting several artworks before leaving the gallery and the goal is to engineer visitors’ experience, (2) each product (artwork) can be displayed at no more than one location, and (3) the visitor’s browsing distribution can depend on the placement of artwork. They show that the problem of optimizing the placement of artwork in the gallery is NP hard. They propose various heuristics to solve the problem and conduct numerical experiments to demonstrate the efficacy of these heuristics but do not analyze the performance of these heuristics theoretically. In our model, the goal is to engineer customer choice and we assume that the browsing distribution is independent of the placement decision but is otherwise arbitrary (and not necessarily Markovian) and we consider a general choice model (not just MNL). We also allow a product to be displayed at multiple locations (in line with retail practice) and focus on developing efficient algorithms with theoretical performance guarantees.
@@
One common approach is to consider a choice model over an expanded universe that consists of all (product, location) pairs. The revenue optimal placement in this model is the optimal solution to an assortment optimization problem with constraints ensuring that every location has at most one product (Abeliuk et al. (2016)). This constraint can be expressed using a partition matroid or a system of totally unimodular (TU) constraints. For MNL choice model, Avadhanula et al. (2016) and Sumida et al. (2021) give polynomial sized LP formulations to solve the TU constrained assortment optimization problem. For the Markov choice model (Blanchet et al. (2016)), Udwani (2025) gave a $\frac{1}{4}$-approximation algorithm for assortment optimization subject to a matroid constraint and Désir et al. (2020) showed that assortment optimization with TU constraints is NP hard to approximate with a factor of $O(n^{1/2-\epsilon})$ for any fixed $\epsilon>0$.
@@
Sequential Search: Our model also bears some resemblance to discrete choice models based on sequential search (Weitzman (1978)). In sequential search, the customer initially does not know the value of the product at any location and searches for the best alternative by examining the locations one by one. There is a cost for examining each location and the search process induces a placement dependent browsing distribution over display locations because the customer’s decision to stop searching may depend on the set of products that they have seen. The setting considered by Derakhshan et al. (2022) is perhaps the most closely related to our work. They consider a product placement problem with sequential search where customers examine the display locations in a fixed order (similar to Placement on a Line) to form a consideration set and then choose at most one product from this set according to a Mixture of MNL choice model. They give novel approximation algorithms and polynomial time approximation schemes for various types of objectives. In contrast, we consider a general store layout but assume that the search process is fully captured by the browsing distribution and independent of product placement. We refer to Wang and Sahin (2018), Kosilova (2022), Ursu et al. (2025) for more detailed reviews of the literature on assortment optimization and discrete choice models based on sequential search.
Assortment Optimization: Recall that the problem of cardinality constrained assortment optimization is a special case of our setting where customers see all products on display. The placement of products does not influence customer choice but the finite number of display locations induces a cardinality constraint. There is a long line of work on assortment optimization for single-purchase choice models. Rusmevichientong et al. (2010) gave a polynomial time algorithm for the problem for MNL choice model. For the Nested Logit choice model, Gallego and Topaloglu (2014) showed that the problem can be solved in polynomial time using linear programming. For the Mixed MNL choice model with a constant number of mixtures, there is a fully polynomial time approximation scheme (FPTAS) (Mittal and Schulz (2013), Désir et al. (2022)). Désir et al. (2020) gave a $0.5-\epsilon$ approximation for the problem (with algorithm runtime proportional to $\frac{1}{\epsilon}$) in the Markov choice model and also showed that the problem is APX-hard. Cardinality constrained assortment optimization has also been studied for various consider-then-choose choice models with placement independent consideration sets. We refer to Aouad et al. (2025b, 2021) for a detailed review of these settings.
@@
Shelf-Space Allocation: There is a significant body of work on the problem of allocating shelf-space in physical stores that sell a variety of (not necessarily substitutable) products (Flamand et al. (2016, 2018, 2023), Hariga et al. (2007)). To the best of our knowledge, the model and methodologies considered in this related literature and our paper are complementary. The common solution methodology in this literature is to formulate a mathematical program (typically a non-linear mixed integer program) that incorporates a variety of important practical considerations in store wide allocation of shelf-space for different product categories that are not substitutable. We refer to Flamand et al. (2023) for a more detailed review of work on shelf-space allocation.